\begin{thebibliography}{}

%\bibitem{An} Y. Andr\'e, Motifs de dimension finie (d'apr\`es S.-I. Kimura, P. O'Sullivan...), S\'eminaire Bourbaki. Vol. 2003/2004. Ast\'erisque 299 Exp. No. 929, viii, 115---145,

%\bibitem{A2} D. Arapura, The Leray spectral sequence is motivic, Invent. math. 160 (2005), 567---589, 

%\bibitem{A} D. Arapura, Varieties with very little transcendental cohomology, in: Motives and algebraic cycles (R. de Jeu and J. Lewis, eds.), Fields Institute Communications, Amer. Math. Soc., Providence 2009,

\bibitem{A2} D. Arapura, Motivation for Hodge cycles, Advances in Math. vol. 207 (2006), 

%\bibitem{AK} D. Arapura and S.-J. Kang, Functoriality of the coniveau filtration, Canad. Math. Bull. 50 (2) (2007), 161---171,

%\bibitem{Ay} J. Ayoub, Motives and algebraic cycles: a selection of conjectures and open questions, preprint,

\bibitem{B0} S. Bloch, On an argument of Mumford in the theory of algebraic cycles, in: G\'eom\'etrie alg\'ebrique (A. Beauville, ed.), Sijthoff and Noordhoff, Angers 1979,

\bibitem{B} S. Bloch, Lectures on algebraic cycles, Duke Univ. Press Durham 1980,

%\bibitem{B2} S. Bloch, Algebraic cycles and higher K--theory, Advances in Math. vol. 61 (1986), 267---304,

%\bibitem{B3} S. Bloch, The moving lemma for higher Chow groups, J. Alg. Geom. 3 (1994), 537---568,

%\bibitem{BO} S. Bloch and A. Ogus, Gersten's conjecture and the homology of schemes, Ann. Sci. Ecole Norm. Sup. 4 (1974), 181---202,

\bibitem{BS} S. Bloch and V. Srinivas, Remarks on correspondences and algebraic cycles, American Journal of Mathematics Vol. 105, No 5 (1983), 1235---1253,

%\bibitem{CM} A. de Cataldo and L. Migliorini, The Hodge theory of algebraic maps, Ann. Scient. Ec. Norm. Sup., 4e s\'erie, t. 38 (2005), 693---750,

%\bibitem{D} A. Durfee, Intersection homology Betti numbers, Proc. Amer. Math. Soc. 123 (1995), 989---993,

%\bibitem{EL} H. Esnault and M. Levine, Surjectivity of cycle maps, in: Journ\'ees de G\'eom\'etrie alg\'ebrique d'Orsay, Ast\'erisque 218 (1993),

%\bibitem{ELV} H. Esnault, M. Levine and E. Viehweg, Chow groups of projective varieties of very small degree, Duke Math. J. 87 (1997), 29---58,

\bibitem{ESV} H. Esnault, V. Srinivas and E. Viehweg, Decomposability of Chow groups implies decomposability of cohomology, in :``Journ\'ees de G\'eom\'etrie Alg\'ebrique d'Orsay, Juillet 1992'', Ast\'erisque 218 (1993), 227---242, 

\bibitem{F} W. Fulton, Intersection theory, Springer--Verlag Berlin Heidelberg New York 1984,

%\bibitem{Fu} B. Fu, Remarks on hard Lefschetz conjectures on Chow groups, Science China Mathematics Vol. 53 No 1 (2010), 105---114,

%\bibitem{FMSS} W. Fulton, R. MacPherson, F. Sottile and B. Sturmfels, Intersection theory on spherical varieties, J. Alg. Geom. 4 (1993), 181---193,

%\bibitem{G} H. Gillet, Riemann--Roch theorems for higher algebraic K--theory, Advances in Math. 40 (1981), 203---289,

%\bibitem{GS} H. Gillet and C. Soul\'e, Descent, motives and K--theory, J. Reine Angew. Math. 478 (1996), 127---176,

%\bibitem{GM} M. Goresky and R. MacPherson, Intersection homology theory, Topology 19 (1980), 135---162,

%\bibitem{GM2} M. Goresky and R. MacPherson, Intersection homology II, Inv. Math. 71 (1983), 77---129,

%\bibitem{GP} V. Guletski\u{\i} and C. Pedrini, The Chow motive of the Godeaux surface, in:
%Algebraic Geometry, a volume in memory of Paolo Francia (M.C. Beltrametti,
%F. Catanese, C. Ciliberto, A. Lanteri and C. Pedrini, editors),
%Walter de Gruyter, Berlin New York, 2002,

%\bibitem{GP2} V. Guletski\u{\i} and C. Pedrini, Finite--dimensional motives and the conjectures of Beilinson and Murre, K--Theory 30(3) (2003), 243---263,

%\bibitem{HS} M. Hanamura and M. Saito, Weight filtrations on the cohomology of algebraic varieties, arxiv 0605603v2,

%\bibitem{H} R. Hartshorne, Equivalence relations on algebraic cycles and subvarieties of small codimension, in: Algebraic geometry, Arcata 1974, Proc. Symp. Pure Math. Vol. 29, Amer. Math. Soc., Providence 1975,


%\bibitem{HI} A. Hirschowitz and J. Iyer, Hilbert schemes of fat $r$--planes and the triviality of Chow groups of complete intersections, in: Vector bundles and complex geometry, Contemp. Math. 522, Amer. Math. Soc., Providence 2010,


%\bibitem{Iv} F. Ivorra, Finite dimensional motives and applications (following S.-I. Kimura, P. O'Sullivan and others), in:
      %  Autour des motifs, Asian-French summer school on algebraic geometry and number theory,
       %Volume III, Panoramas et synth\`eses, Soci\'et\'e math\'ematique de France 2011,

%\bibitem{I} J. Iyer, Murre's conjectures and explicit Chow--K\"unneth projectors for varieties with a nef tangent bundle, Transactions of the Amer. Math. Soc. 361 (2008), 1667---1681,

%\bibitem{J} U. Jannsen, Mixed motives and algebraic K--theory, Springer Lecture Notes in Mathematics 1400 (1990),

%\bibitem{J1} U. Jannsen, 
%Motives, numerical equivalence, and semi-simplicity, Invent. Math. 107(3) (1992), 447---452, 

%\bibitem{J2} U. Jannsen, Motivic sheaves and filtrations on Chow groups, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1,  

%\bibitem{J3} U. Jannsen, Equivalence relations on algebraic cycles, in: The arithmetic and geometry
%of algebraic cycles (B. Gordon et alii, eds.), Banff Conf. 1998, Kluwer,

%\bibitem{J4} U. Jannsen, On finite--dimensional motives and Murre's conjecture, in: Algebraic cycles and motives, volume 344 of London Math. Soc. Lecture Note Ser., Cambridge Univ. Press, Cambridge 2007,

%\bibitem{Kah} B. Kahn, Equivalences rationnelle et num\'erique sur certaines vari\'et\'es de type ab\'elien sur un corps fini, Ann. Sci. Ec. Norm. Sup. 36 (2003), 977---1002,

\bibitem{KMP} B. Kahn, J. P. Murre and C. Pedrini, On the transcendental part of the motive of a surface, in: Algebraic cycles and motives. Vol. 2, volume 344 of London Math. Soc. Lecture Note Ser., 143---202, Cambridge Univ. Press, Cambridge 2007,


%\bibitem{KS} B. Kahn and R. Sujatha, Birational motives, I: pure birational motives, arXiv:0902.4902v4,

%\bibitem{Kim} S. Kimura, Fractional intersection and bivariant theory, Comm. in Algebra 20 (1992), 285---302, 

%\bibitem{Kim} S. Kimura, Chow groups are finite dimensional, in some sense,
%Math. Ann. 331 (2005), 173---201,

%\bibitem{Kim2} S. Kimura, Surjectivity of the cycle map for Chow motives, in: Motives and algebraic cycles, Fields Institute Comm. vol. 56, Amer. Math. Soc., Providence 2009,

\bibitem{K0} S. Kleiman, Algebraic cycles and the Weil conjectures, in: Dix expos\'es sur la cohomologie des sch\'emas, North--Holland Amsterdam, 1968, 359---386, 

\bibitem{K} S. Kleiman, The standard conjectures, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1, 

\bibitem{yet} R. Laterveer, Yet another version of Mumford's theorem, Archiv Math. 104 (2015), 125---131, 

\bibitem{moi} R. Laterveer, On a multiplicative version of Bloch's conjecture, submitted to Beitr\"age zur Algebra und Geometrie,


%\bibitem{Lat} R. Laterveer, A brief note concerning hard Lefschetz for Chow groups, to appear in Canadian Math. Bulletin, http://dx.doi.org/10.4153/CMB-2015-046-x,

%\bibitem{moi} R. Laterveer, Algebraic varieties with small Chow groups, J. Math. Kyoto Univ. Vol. 38 No 4 (1998), 673---694,

%\bibitem{Lev} M. Levine, Techniques of localization in the theory of algebraic cycles, J. Alg. Geom. 10
%(2001), 299---363,

%\bibitem{L} J. Lewis, Towards a generalization of Mumford's theorem, J. Math. Kyoto Univ. 29 (1989), 195---204,

%\bibitem{L2} J. Lewis, A generalization of Mumford's theorem, II, Illinois Journal of Mathematics, Vol. 39 No 2 (1995), 288---304,

%\bibitem{LH} J. Lewis, A survey of the Hodge conjecture, Second edition, CRM Monograph series, Amer. Math. Soc., Providence 1999,


\bibitem{M} D. Mumford, Rational equivalence of $0$--cycles on surfaces, J. Math. Kyoto Univ. Vol. 9 No 2 (1969), 195---204,

\bibitem{Mu} J. Murre, On a conjectural filtration on the Chow groups of an algebraic variety, Indag. Math. (N.S.) 4 (1993), 177---201,

\bibitem{MNP} J. Murre, J. Nagel and C. Peters, Lectures on the theory of pure motives, Amer. Math. Soc. University Lecture Series 61, 2013,

%\bibitem{P} C. Pedrini, On the finite dimensionality of a $K3$ surface, Manuscripta Mathematica 138 (2012), 59---72,

%\bibitem{PS} C. Peters and J. Steenbrink, Mixed Hodge structures, Springer Verlag,

%\bibitem{Q} D. Quillen, Higher algebraic K--theory: I, in: Algebraic K-theory, I: Higher K-theories (Proc. Conf., Battelle Memorial Inst., Seattle, Wash., 1972), Lecture Notes in Math 341, Springer-Verlag Berlin New York,

%\bibitem{R} A.A. Rojtman, The torsion of the group of 0--cycles modulo rational equivalence, Annals of Mathematics 111 (1980), 553---569,

%\bibitem{S} C. Schoen, On Hodge structures and non--representability of Chow groups, Comp. Math. 88 (1993), 285---316,

\bibitem{Sch} A. Scholl, Classical motives, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1,  

\bibitem{Tan} S. Tankeev, On the standard conjecture of Lefschetz type for complex projective threefolds. II, Izvestiya Math. 75:5 (2011), 1047---1062,

%\bibitem{T} B. Totaro, Chow groups, Chow cohomology, and linear varieties, Forum of Mathematics, Sigma (2014), vol. 1, e1,

\bibitem{V} C. Vial, Algebraic cycles and fibrations, Documenta Math. 18 (2013), 1521---1553,

\bibitem{V2} C. Vial, Projectors on the intermediate algebraic Jacobians, New York J. Math. 19 (2013), 793---822,

%\bibitem{V3} C. Vial, Remarks on motives of abelian type, arxiv:1112.1080v2,

%\bibitem{V4} C. Vial, Niveau and coniveau filtrations on cohomology groups and Chow groups, Proceedings of the LMS 106(2) (2013), 410---444,

%\bibitem{Vial} C. Vial, Chow-K\"unneth decomposition for 3- and 4-folds fibred by varieties with trivial Chow group of zero-cycles, J. Algebraic Geom. 24 (2015), 51---80,

%\bibitem{Voe} V. Voevodsky, A nilpotence theorem for cycles algebraically equivalent to zero, Internat. Math. Research Notices 4 (1995), 187---198,

%\bibitem{V9} C. Voisin, Remarks on zero--cycles on self--products of varieties, in: Moduli of vector bundles (Proceedings of the Taniguchi Congress), Maruyama Ed., Decker 1994,

%\bibitem{V0} C. Voisin, The generalized Hodge and Bloch conjectures are equivalent for general complete intersections, Annales scientifiques de l'ENS 46, fascicule 3 (2013), 449---475,

%\bibitem{V1} C. Voisin, The generalized Hodge and Bloch conjectures are equivalent for general complete intersections, II, \`a para\^itre au Journal of Mathematical Sciences, the University of Tokyo, Kodaira Centennial issue,

\bibitem{Vo} C. Voisin, Chow Rings, Decomposition of the Diagonal, and the Topology of Families, Princeton University Press, Princeton and Oxford, 2014,

%\bibitem{Y} Q. Yin, Finite--dimensionality and cycles on powers of K3 surfaces, Comment. Math. Helv. 90 (2015), 503---511.

%\bibitem{Z} Q. Zhang, Rational connectedness of log $\QQ$--Fano varieties, J. Reine Angew. Math. 590 (2006), 131---142. 



\end{thebibliography}

\RequirePackage{fix-cm}
\documentclass[smallcondensed]{svjour3}     
\smartqed  
\usepackage{graphicx}
\usepackage{times,a4wide,mathrsfs}
\usepackage{amsmath}
\usepackage{amsfonts}

\newcommand{\C}{\mathbb{C}}

\newcommand{\QQ}{\mathbb{Q}}
\newcommand{\NN}{\mathbb{N}}

\newcommand{\OO}{\mathcal O}

\newcommand{\id}{\hbox{id}}

\newcommand{\gr}{\hbox{Gr}}
\newcommand{\wt}{\widetilde}
\newcommand{\ima}{\hbox{Im}}

\newtheorem{convention}{Conventions}

\newtheorem{nonumbering}{Theorem}

\begin{document}

\title{On a multiplicative version of Mumford's theorem
}


\author{Robert Laterveer 
}


\institute{CNRS - IRMA, Universit\'e de Strasbourg \at
              7 rue Ren\'e Descartes \\
              67084 Strasbourg cedex\\
              France\\
              \email{laterv@math.unistra.fr}           
           }

\date{Received: date / Accepted: date}

\maketitle

\begin{abstract} A theorem of Esnault, Srinivas and Viehweg asserts that if the Chow group of $0$--cycles of a smooth complete complex variety decomposes, then the top--degree coherent cohomology group decomposes similarly. In this note, we prove a similar statement for Chow groups of arbitrary codimension, provided the variety satisfies the Lefschetz standard conjecture.

\end{abstract}

\keywords{Algebraic cycles \and Chow groups \and Mumford's theorem \and Intersection product}
 \subclass{ 14C15 \and  14C25 \and  14C30}

\section{Introduction}

Since Mumford's famous 1969 paper \cite{M}, it is well--known that the Chow group of $0$--cycles $A^nX$ on a complex variety $X$ influences the cohomology group $H^n(X,\QQ)$:

\begin{nonumbering}[Mumford \cite{M}] Let $X$ be a smooth complete variety of dimension $n$ defined over $\C$. Suppose that $A^nX_{\QQ}$ is supported on a divisor. Then $H^n(X,\QQ)$ is supported on a divisor, in particular $H^n(X,\OO_X)=0$.
\end{nonumbering}

In the 1992 paper \cite{ESV}, Esnault, Srinivas and Viehweg study the multiplicative behaviour of the Chow ring $A^\ast X$ versus the multiplicative behaviour of various cohomology rings associated to $X$. We now state the part of their result that is relevant to us. For a given partition 
   $n=n_1+\cdots+n_r$
   (with $n_i\in\NN_{>0}$),
let us consider the following properties:

\vskip0.5cm
\noindent
(P1) There exists a Zariski open $V\subset X$, such that intersection product induces a surjection
  \[   A^{n_1}V_{\QQ}\otimes A^{n_2}V_{\QQ}\otimes\cdots\otimes A^{n_r}V_{\QQ}\ \to\ A^nV_{\QQ}\ ;\]

  \noindent
  (P2) There exists a Zariski open $V\subset X$, such that cup product induces a surjection
  \[ H^{n_1}(V,\QQ)\otimes H^{n_2}(V,\QQ)\otimes\cdots\otimes H^{n_r}(V,\QQ)\ \to\ H^n(V,\QQ)/N^1\ \]
  (here $N^\ast$ denotes the coniveau filtration);
  
  \noindent
  (P3) Cup product induces a surjection
    \[ H^{n_1}(X,\OO_X)\otimes H^{n_2}(X,\OO_X)\otimes\cdots\otimes H^{n_r}(X,\OO_X)\ \to\ H^n(X,\OO_X)\ .\]

\vskip0.5cm

In these terms, what Esnault, Srinivas and Viehweg prove is the following:

\begin{nonumbering}[Esnault--Srinivas--Viehweg \cite{ESV}] Let $X$ be a smooth complete variety of dimension $n$ over $\C$. Then (P1) implies (P3), and (P2) implies (P3).
\end{nonumbering}

The implication from (P1) to (P3) is a kind of multiplicative variant of Mumford's theorem, and the proof in \cite{ESV} is motivated by Bloch's proof of Mumford's theorem using a ``decomposition of the diagonal'' argument (\cite{B}, \cite{B0}, cf. also \cite{BS}). As noted in \cite[remark 2]{ESV}, the generalized Hodge conjecture would imply that (P2) and (P3) are equivalent.\footnote{It is somewhat frustrating that it is not known unconditionally whether (P1) implies (P2), i.e. without assuming the generalized Hodge conjecture. Apparently Esnault, Srinivas and Viehweg had claimed to prove this in an earlier version of their paper, but the argument was found to be incomplete \cite[remark 2]{ESV}.}

In this note, we show that the Esnault--Srinivas--Viehweg theorem can be extended from $0$--cycles to arbitrary Chow groups. This is possible provided the variety $X$ satisfies the Lefschetz standard conjecture $B(X)$ (this is analogous to \cite{yet}, where I extended Mumford's theorem from $0$--cycles to arbitrary Chow groups, provided $B(X)$ holds):

\begin{nonumbering}[(=theorem \ref{main})] Let $X$ be a smooth projective variety of dimension $n$ over $\C$ that satisfies $B(X)$. Suppose there exists a Zariski open $V\subset X$, and $j=j_1+\cdots+ j_r$ with $j_i\in\NN_{>0}$ such that intersection product induces a surjection
  \[  A^{j_1}V_{\QQ}\otimes A^{j_2}V_{\QQ}\otimes\cdots\otimes A^{j_r}V_{\QQ}\ \to\ A^jV_{\QQ}\ .\]
Then cup product induces a surjection
  \[ H^{j_1}(X,\OO_X)\otimes H^{j_2}(X,\OO_X)\otimes\cdots\otimes H^{j_r}(X,\OO_X)\ \to\ H^j(X,\OO_X)\ .\]
\end{nonumbering}

The proof of this theorem, which is very similar to the proof given by Esnault--Srinivas--Viehweg in \cite{ESV}, is an exercise in using the meccano of correspondences and the Bloch--Srinivas formalism.

It seems natural to wonder whether the converse to theorem \ref{main} might perhaps be true (this would be a multiplicative variant of Bloch's conjecture). In \cite{moi}, I prove this converse implication in some special cases for $0$--cycles (i.e. $j=n$); the converse implication for $j\not=n$ appears to be more difficult.

\begin{convention} In this note, the word {\sl variety\/} will refer to a quasi--projective irreducible algebraic variety over $\C$, endowed with the Zariski topology. A {\sl subvariety\/} is a (possibly reducible) reduced subscheme which is equidimensional. The Chow group of $j$--dimensional algebraic cycles on $X$ with $\QQ$--coefficients modulo rational equivalence is denoted $A_jX$; for $X$ smooth of dimension $n$ the notations $A_jX$ and $A^{n-j}X$ will be used interchangeably. Caveat: note that what we denote $A^jX$ is elsewhere often denoted $CH^j(X)_{\QQ}$.
In an effort to lighten notation, we will often write $H^jX$ or $H_jX$ to indicate singular cohomology $H^j(X,\QQ)$ resp. Borel--Moore homology $H_j(X,\QQ)$.

For basics concerning algebraic cycles and their functorial behaviour, the curious reader is invited to consult \cite{F}. For the formalism of correspondences, cf. \cite{Sch}, \cite{MNP}.
\end{convention}

\section{Preliminary}

Let $X$ be a smooth projective variety of dimension $n$, and $h\in H^2(X,\QQ)$ the class of an ample line bundle. The hard Lefschetz theorem asserts that the map
  \[  L^{n-i}\colon H^i(X,\QQ)\to H^{2n-i}(X,\QQ)\]
  obtained by cupping with $h^{n-i}$ is an isomorphism, for any $i< n$. One of the standard conjectures asserts that the inverse isomorphism is algebraic.

\begin{definition}[Lefschetz standard conjecture]{} Given a variety $X$, we say that $B(X)$ holds if for all ample $h$, and all $i<n$ the isomorphism 
  \[  (L^{n-i})^{-1}\colon 
  H^{2n-i}(X,\QQ)\stackrel{\cong}{\rightarrow} H^i(X,\QQ)\]
  is induced by a correspondence.
 \end{definition}  
 
 \begin{remark} It is known that $B(X)$ holds for the following varieties: curves, surfaces, abelian varieties \cite{K0}, \cite{K}, threefolds not of general type \cite{Tan}, varieties motivated by a surface in the sense of Arapura \cite{A2} (this includes the Hilbert schemes of $0$--dimensional subschemes of surfaces \cite[Corollary 7.5]{A2}), 
 $n$--dimensional varieties $X$ which have $A_i(X)_{}$ supported on a subvariety of dimension $i+2$ for all $i\le{n-3\over 2}$ \cite[Theorem 7.1]{V}, $n$--dimensional varieties $X$ which have $H_i(X)=N^{\llcorner {i\over 2}\lrcorner}H_i(X)$ for all $i>n$ \cite[Theorem 4.2]{V2}, products and hyperplane sections of any of these \cite{K0}, \cite{K}.
 \end{remark}

It is known that $B(X)$ implies that the K\"unneth components
  \[  \pi_j\in H^{2n-j}(X)\otimes H^j(X)\subset H^{2n}(X\times X)\]
  of the diagonal $\Delta\subset X\times X$ are algebraic \cite{K0}, \cite{K}. Moreover, these K\"unneth components satisfy the following property:
  
  \begin{lemma}\label{kunneth} Let $X$ be a smooth projective variety satisfying $B(X)$, and let $h\in H^2(X)$ be the class of an ample line bundle.
  For any $j\le n$, there exists a cycle $P_j\in A^j(X\times X)$ such that
    \[   \pi_j=(\tau\times\id)_\ast(\tau\times\id)^\ast(P_j)\ \ \in H^{2n}(X\times X)\ ,\]
    where $\tau\colon Y_j\to X$ denotes the inclusion of a dimension $j$ complete intersection of class $[Y_j]=h^{n-j}$.
  \end{lemma}

\begin{proof} As mentioned above, $B(X)$ ensures that $\pi_j$ is algebraic \cite{K0}, \cite{K}. Consider now the isomorphism
  \[   L^{n-j}\times\id\colon\ \ H^jX\otimes H^jX\ \xrightarrow{\cong}\ H^{2n-j}X\otimes H^jX\ \]
  (here we tacitly identify both sides with their images in $H^\ast(X\times X)$).
  
  Since we have $B(X)$, there exists a correspondence, say $Q\in A^{j}(X\times X)$, such that
    \[  (L^{n-j}\times\id)   (Q\times\id)_\ast=\id\colon\ \ H^{2n-j}X\otimes H^jX\ \to\ H^{2n-j}X\otimes H^jX\ .\]
    Since $\pi_j$ is algebraic,
      \[ P_j:= (Q\times\id)_\ast (\pi_j)\in A^j(X\times X)\]
      is still algebraic, and has the requested property.
\end{proof}

\begin{remark} Lemma \ref{kunneth} implies in particular that for a variety satisfying $B(X)$, the K\"unneth component $\pi_j$ is represented by an algebraic cycle contained in $Y_j\times X$, for a dimension $j$ complete intersection $Y_j$. This was also proven in \cite{KMP} (and independently in \cite[proof of theorem 3.1]{yet}, as I wasn't aware of the Kahn--Murre--Pedrini reference at the time).
\end{remark}

\section{Main}

We now prove the main theorem of this note:

\begin{theorem}\label{main} Let $X$ be a smooth projective variety of dimension $n$ over $\C$ that satisfies $B(X)$. Suppose there exists a Zariski open $V\subset X$, and $j=j_1+\cdots+ j_r$ with $j_i\in\NN_{>0}$ such that intersection product induces a surjection
  \[  A^{j_1}V_{}\otimes A^{j_2}V_{}\otimes\cdots\otimes A^{j_r}V_{}\ \to\ A^jV_{}\ .\]
Then cup product induces a surjection
  \[ H^{j_1}(X,\OO_X)\otimes H^{j_2}(X,\OO_X)\otimes\cdots\otimes H^{j_r}(X,\OO_X)\ \to\ H^j(X,\OO_X)\ .\]
\end{theorem}

\begin{proof} Since $B(X)$ holds, it follows from lemma \ref{kunneth}
that the K\"unneth component $\pi_j$ can be written
    \[   \pi_j=(\tau\times\id)_\ast(\tau\times\id)^\ast(P_j)\ \ \in H^{2n}(X\times X)\ \]
    for some $P_j\in A^j(X\times X)$,
    where $\tau\colon Y_j\to X$ is the inclusion of a dimension $j$ complete intersection.
    
 Applying the Bloch--Srinivas argument, in the form of proposition \ref{BS} below, to the cycle $P_j\in A^j(X\times X)$, we find a decomposition
   \[   P_j=C_1\cdot \ldots \cdot C_r+\Gamma_1+\Gamma_2\ \ \in A^j(X\times X)\ ,\]
   where $\Gamma_1, \Gamma_2$ are supported on $D\times X$, resp. on $X\times D$, for some divisor $D\subset X$. This induces a decomposition of the K\"unneth component
   \[  \begin{split}\pi_j&=(\tau\times\id)_\ast(\tau\times\id)^\ast(C_1\cdot \ldots \cdot C_r)+
     (\tau\times\id)_\ast(\tau\times\id)^\ast   (\Gamma_1+\Gamma_2)\\
                                 &=(\tau\times\id)_\ast(\tau\times\id)^\ast(C_1\cdot \ldots \cdot C_r)+
        \Gamma_1^\prime+\Gamma_2^\prime      \ \ \in H^{2n}(X\times X)\ ,\end{split}\]
 where $\Gamma_2^\prime$ is still supported on $X\times D$, and $\Gamma_1^\prime$ is supported on $Z\times X$, for some $Z\subset X$ of dimension $j-1$ (indeed, the general complete intersection $Y_j$ will be in general position with respect to $D$; we then define $Z$ to be $D\cap Y_j$).
 
 Now we consider the action of $\pi_j$ on $H^j(X,\OO_X)$. Since $H^j(X,\OO_X)=\gr^0_F H^j(X,\C)$ (where $F$ is the Hodge filtration), $\pi_j$ acts as the identity on $H^j(X,\OO_X)$. On the other hand, it is clear that
   \[ (\Gamma_1^\prime)_\ast H^j(X,\OO_X)=0\]
   (by lemma \ref{comp2}, the action of $\Gamma_1^\prime$ factors over $\gr^0_F H^j(Z,\C)$, which is $0$ for dimension reasons), and also that
   \[ (\Gamma_2^\prime)_\ast H^j(X,\OO_X)=0\]
   (by lemma \ref{comp2}, the action of $\Gamma_2$ factors over $\gr^{-1}_F H^{j-2}(\wt{D},\C)=0$, where $\wt{D}$ is a resolution of singularities of $D$).
   To finish the argument, it only remains to analyze the action
   \[   \bigl((\tau\times\id)_\ast(\tau\times\id)^\ast(C_1\cdot \ldots \cdot C_r)\bigr)_\ast\colon\ \ H^j(X,\OO_X)\ \to\ H^j(X,\OO_X)\ .\]
   
   Using lemmas \ref{comp} and \ref{comp2}, we find an inclusion
   \[   \bigl((\tau\times\id)_\ast(\tau\times\id)^\ast(C_1\cdot \ldots \cdot C_r)\bigr)_\ast H^j(X,\OO_X)     \ \subset\ \bigl(C_1\cdot \ldots \cdot C_r\bigr)_\ast \gr_F^{n-j} H^{2n-j}(X,\C)\ .\]
   
   Using lemma \ref{esv}, we find that
   \[   \bigl(C_1\cdot \ldots \cdot C_r\bigr)_\ast \gr_F^{n-j} H^{2n-j}(X,\C)\subset  \ima\bigl(H^{j_1}(X,\OO_X)\otimes\cdots\otimes H^{j_r}(X,\OO_X)\ \to\ H^j(X,\OO_X)\bigr) \ ,\]
   and so we are done.
  \end{proof}

 \begin{proposition}[Bloch--Srinivas style]\label{BS} Let $X$ be a smooth projective variety of dimension $n$. Suppose there exists a Zariski open $V\subset X$, and $j=j_1+\cdots+ j_r$ with $j_i\in\NN_{>0}$ such that intersection product induces a surjection
  \[  A^{j_1}V_{}\otimes A^{j_2}V_{}\otimes\cdots\otimes A^{j_r}V_{}\ \to\ A^jV_{}\ .\]
  Then for any $a\in A^j(X\times X)$, there exists a decomposition
  \[   a= C_1\cdot\ldots\cdot C_r+\Gamma_1+\Gamma_2\ \ \in A^j(X\times X)\ ,\]
  where $C_i\in A^{j_i}(X\times X)$, and $\Gamma_1, \Gamma_2$ are supported on $D\times X$ (resp. on $X\times D$), for some divisor $D\subset X$.
  \end{proposition}
  
  \begin{proof} To be sure, this is a variant of the argument of \cite{BS}, exploiting the fact that $\C$ is a universal domain.
  Let $D_1\subset X$ denote the complement of $V$. Taking the smallest possible field of definition, we can suppose everything ($X$, $V$ and the cycle $a$) is defined over a field $k\subset\C$ which is finitely generated over its prime subfield. Then the inclusion $k(X)\subset\C$ (which comes from $\C$ being a universal domain) induces an injection
  \[  A^j(X_{k(X)})\ \to\ A^j(X_{\C})\ \]
  \cite[Appendix to Lecture 1]{B}. On the other hand,
  \[ A^j(X_{k(X)})_{}=   \varinjlim A^j(X\times U)_{}\ ,\]
 where the limit is taken over opens $U\subset X$ \cite[Appendix to Lecture 1]{B}.   
 
 Given the cycle $a\in A^j(X\times X)$, consider the restriction
   \[ a_{restr}\ \ \in A^j(X_{k(X)})\ .\]
   The assumption implies there exist cycles $c_i\in A^{j_i}(X_{\C})$ such that
   \[  a_{restr}=c_1\cdot\ldots\cdot c_r+a_0\ \ \in A^j(X_{\C})\ ,\]
   where $a_0$ is supported on $D_1$. Now, we extend $k$ so that the cycles $c_i$ are also defined over $k$ (and $k$ is still finitely generated over its prime subfield, so that $k(X)\subset\C$). Then using the injection $A^j(X_{k(X)})\ \to\ A^j(X_{\C})$ cited above, we obtain the decomposition
   \[  a_{restr}=c_1\cdot\ldots\cdot c_r+a_0\ \ \in A^j(X_{k(X)})\ .\]  
   Let $C_i\in A^{j_i}(X\times X)$ be any cycle restricting to $c_i$, and let $\Gamma_1$ be any cycle restricting to $a_0$. Then using the limit property cited above, we find that the difference
   \[ a-C_1\cdot\ldots\cdot C_r-\Gamma_1   \ \ \in A^j(X\times X)\]
   restricts to $0$ in $A^j(X\times U)$, for some open $U\subset X$. This means there exists a divisor $D_2\subset X$ and a cycle $\Gamma_2$ supported on $D_2$ such that
   \[   a= C_1\cdot\ldots\cdot C_r+\Gamma_1+\Gamma_2\ \ \in A^j(X\times X)\ .\]
   Taking $D$ a divisor containing both $D_1$ and $D_2$, this proves the proposition.
   \end{proof}
  
  \begin{lemma}\label{comp} Let $f\colon Y\to X$ be a proper morphism of smooth projective varieties, where $\dim X=n$ and $\dim Y=m$. Let $C\in A^j(X\times X)$.
  Then
    \[  \bigl( (f\times\id)^\ast C\bigr)_\ast = C_\ast f_\ast\colon\ \ H^iY\ \to\ H^{i+2(j-m)}X\ .\]
     \end{lemma}
  
  \begin{proof} This is purely formal, and surely well--known. Let $p_1, p_2\colon X\times X\to X$ denote projection on the first (resp. second) factor. Let $q_1, q_2$ denote projections from $Y\times X$ to $Y$ (resp. to $X$).
  For $a\in H^iY$, we have
    \[\begin{split} \bigl( (f\times\id)^\ast C\bigr)_\ast(a)&=   (q_2)_\ast \bigl( (q_1)^\ast(a)\cdot (f\times\id)^\ast (C)\bigr)\\
                    &= (p_2)_\ast \bigl( (f\times\id)_\ast ((q_1)^\ast(a)\cdot (f\times\id)^\ast(C))  \bigr) \\
                    &=(p_2)_\ast \bigl( (f\times\id)_\ast (q_1)^\ast(a)\cdot C\bigr)\\
                    &=(p_2)_\ast \bigl( (p_1)^\ast f_\ast (a)\cdot C\bigr)=: C_\ast f_\ast(a)\ \ \in H^{i+2(j-m)}X\ .\\
                    \end{split}\]
    \end{proof}

\begin{lemma}\label{comp2} Let $f\colon Y\to X$ be as in lemma \ref{comp}. Let $D\in A^j(Y\times X)$. Then
  \[   \bigl( (f\times\id)_\ast D\bigr)_\ast = D_\ast f^\ast\colon\ \ H^iX\ \to\ H^{i+2(j-m)}X\ .\]
  \end{lemma}
  
  \begin{proof} Just as lemma \ref{comp}, this is surely well--known. Keeping the notation of lemma \ref{comp}, for $b\in H^iX$ we have
    \[ \begin{split} \bigl( (f\times\id)_\ast D\bigr)_\ast (b)  &= (p_2)_\ast \bigl( (p_1)^\ast(b)\cdot (f\times\id)_\ast (D)\bigr)\\
                                                      &= (p_2)_\ast \bigl(   (f\times\id)_\ast \bigl( (f\times\id)^\ast (p_1)^\ast(b)\cdot D\bigr)\bigr)\\
                                                      &=(p_2)_\ast (f\times\id)_\ast \bigl(  (q_1)^\ast f^\ast(b)\cdot D\bigr)\\
                                                      &= (q_2)_\ast \bigl(  (q_1)^\ast f^\ast(b)\cdot D\bigr)=: D_\ast f^\ast(b)\ \ \in H^{i+2(j-m)}X\ .  \\
                                                         \end{split}\]
  \end{proof}

  \begin{lemma}[Esnault--Srinivas--Viehweg \cite{ESV}]\label{esv} Let $X$ be a smooth projective variety of dimension $n$. Let $C_i\in A^{j_i}(X\times X), i=1,\ldots,r$, with $j=j_1+\cdots+j_r$. Then
    \[   (C_1\cdot\ldots\cdot C_r)_\ast \gr_F^{n-j}H^{2n-j}(X,\C)\subset\ima\Bigl( H^{j_1}(X,\OO_X)\otimes\cdots\otimes H^{j_r}(X,\OO_X)\ \to\ H^j(X,\OO_X)\Bigr)\ .\]  
  \end{lemma}
  
  \begin{proof} This is shamelessly plagiarized from \cite{ESV}, who prove the $j=n$ case. Let
  $ C\in A^j(X\times X)$
  be any correspondence.
    The crucial observation is that the action
    \[  C_\ast\colon\ \ \gr_F^{n-j}H^{2n-j}(X,\C)\ \to\ \gr_F^0 H^j(X,\C)\]
  only depends on the image of $C$ under the composite map
  \[ \iota\colon\ \   A^j(X\times X)\ \to\  H^{2j}(X\times X,\C)\ \to\ H^jX\otimes H^jX\ \to\ \gr_F^j H^jX\otimes \gr_F^0 H^jX\ \]
  (Here the second map is given by the K\"unneth decomposition, and the last map is induced by projection on the appropriate summands of the Hodge decomposition).
  Indeed, suppose $C\in A^j(X\times X)$ is such that $\iota(C)=0$, i.e. the K\"unneth part of type $H^jX\otimes H^jX$ of $C$ is contained in 
     \[ \bigoplus_{i<j}  \gr_F^i H^jX\otimes \gr_F^{j-i} H^jX\ \subset\  \gr_F^j (H^jX\otimes H^jX)\ .\]
     Then, for $a\in \gr_F^{n-j}H^{2n-j}(X,\C)$ we find that 
     \[  (p_1)^\ast(a)\cdot C\ \ \in \bigoplus_{i<j} \gr_F^{n-j+i} H^{2n}X\otimes \gr_F^{j-i}H^jX=0\ ,\]
     and hence
     \[  C_\ast (a)=0\ \ \in H^j(X,\C)\ .\]
   Next, we apply this observation to
     \[ C=C_1\cdot\ldots\cdot C_r\ \ \in A^j(X\times X)\ ,\]
     with $C_i\in A^{j_i}X$.     
      The Hodge decomposition then gives that
     \[  \begin{split}\iota(C)=\iota(C_1)\cdot\ldots\cdot\iota(C_r)\in\ &\ima\bigl(\gr_F^{j_1} H^{j_1}X\otimes\cdots\otimes \gr_F^{j_r} H^{j_r}X\bigr)\otimes  \ima\bigl(\gr_F^{0} H^{j_1}X\otimes\cdots\otimes \gr_F^{0} H^{j_r}X\bigr)\\
                          &  \ \subset\    \gr_F^j H^jX\otimes \gr_F^0 H^jX\ .\\
                          \end{split}\]
     This proves the lemma: suppose
     \[ \iota(C)=\sum_k C^k_{left}\otimes C^k_{right}\ \ \in \gr_F^j H^jX\otimes \gr_F^0 H^jX\ .\]
     Then reasoning as above, we find that
     \[  (\iota(C))_\ast(a)= \sum_k (p_2)_\ast \bigl( (a\cup C^k_{left})\otimes C^k_{right} \bigr)=\sum_k \alpha_k C^k_{right}\ \ \in \gr_F^0 H^jX\ ,\]
     where the $\alpha_k$ are complex numbers (this is because $H^{2n}X$ is one--dimensional and generated by the class of a point). 
     \end{proof}

\begin{remark} It is mainly the contrapositive of theorem \ref{main} that is useful (this is another remark made in \cite{ESV} for their theorem). Indeed, suppose $X$ and $j=j_1+\ldots+j_r$ are such that
  \[   H^{j_1}(X,\OO_X)\otimes H^{j_2}(X,\OO_X)\otimes\cdots\otimes H^{j_r}(X,\OO_X)\ \to\ H^j(X,\OO_X)\ \]
is {\em not\/} surjective (for example, because
  \[ \prod_{i=1}^r \dim H^{j_i}(X,\OO_X)< \dim H^j(X,\OO_X)\ ).\]  
  Then by theorem \ref{main}, likewise
  \[     A^{j_1}X_{}\otimes \cdots\otimes A^{j_r}X_{}\ \to\ A^jX_{}\ \]
  fails to be surjective (and the same holds for any open $V\subset X$).
  \end{remark}

\begin{acknowledgements}
This note is the fruit of the Strasbourg 2014---2015 ``groupe de travail'' based on the monograph \cite{Vo}. I want to thank all the participants of this groupe de travail for the very pleasant and stimulating atmosphere, and their interesting lectures. Furthermore, many thanks to Yasuyo, Kai and Len for providing a wonderful working environment at home in Schiltigheim. 
Special thanks to Kai for all the bicycle trips made together this summer.
\end{acknowledgements}



\end{document}